# Vision-LSTM: xLSTM as Generic Vision Backbone

### v1

@@ A.1 Runtime Comparison of ViL vs Vim
Optimization (T)iny (S)mall (B)ase (L)arge
@@
A.2 Impact of Longer Training
We investigate training for a longer duration on ViL-T+ and train it for 400 and 800 epochs following the hyperparameters in Table 5. After 400 epochs, ViL-T+ achieves 77.2% accuracy and after doubling the training duration the model achieves 78.1% accuracy.
Our implementation of DeiT-III-T achieves 75.6% after 400 epochs and 76.2% with 800 epochs using the hyperparameters in Table 6.
Appendix B Implementation Details
B.1 Hardware
@@
We estimate the total number of A100 GPU-hours used for this project to be 25K hours. This estimate includes everything from initial exploration, method development, analysis and evaluations.
B.2 FLOPS Calculation
We calculate the FLOPS from Table 2 with the fvcore library [^1] https://github.com/facebookresearch/fvcore. As fvcore does not support all operations of ViL, the FLOPS are not 100% accurate, but nevertheless serve as a good reverence point to compare the relative computational effort between block designs.
B.3 ViL Hyperparameters
Table 5 shows detailed hyperparameters used to train ViL models. In the “long-sequence fine-tuning” setting, we follow Vim [44] and fine-tune for 30 epochs with a learning rate of 1e-5, a weight decay of 1e-8 using an overlap of 50% (8 pixels) between consecutive patches. This increases the sequence length from 196 to 729. Other hyperparameters follow Table 5.
@@ B.3 ViL Hyperparameters
Epochs 800 (T), 400 (S/B)
Batch size 1024 (S), 2048 (T/B)
@@
   Base Learning rate 1e-3
@@
Learning rate schedule linear warmup $\rightarrow$ cosine decay
   Warmup epochs 5
@@
Table 5: Hyperparameters for training ViL on ImageNet-1K, inspired by DeiT-III [37].
B.4 DeiT-III Reimplementation Hyperparameters
@@
Batch size 2048
@@
   Base Learning rate 1e-3
@@
Learning rate schedule linear warmup $\rightarrow$ cosine decay

### v2

@@ B.1 Runtime Comparison of ViL vs Vim
Model Optimization (T)iny (S)mall (B)ase (L)arge
@@
B.2 Impact of Longer Training
We investigate the impact of training for a longer duration in Table 8. We also attempted to train larger models for longer but did not observe similar performance gains. One could probably change that by tuning hyperparameters for an 800 epoch training schedule, which we omit due to the large computational costs associated with training larger models.
Model Epochs IN-1K ACC VTAB-1K ADE20K mIoU
DeiT-III-T 400 75.6 67.0 39.1
DeiT-III-T 800 76.2 67.1 39.8
ViL-T 400 77.2 67.8 40.9
ViL-T 800 78.3 68.3 41.2
Table 8: Performance comparison of tiny models when training for 400 and 800 epochs. ADE20K mIoU uses single-scale evaluation. All settings follow the ones used in the main paper.
B.3 Long-sequence fine-tuning
Table 9 shows results for the long-sequence fine-tuning setting of Vim [67].
Method #Params ACC
DeiT-III-T (reimpl.) 6M 78.3
Vim-T [67] 7M 78.3
ViL-T 6M 80.1
DeiT-S [51] 22M 81.4
DeiT-III-S [54] 22M 83.2
Vim-S [67] 26M 81.6
ViL-S 23M 82.7
DeiT-B [51] 86M 83.2
DeiT-III-B [54] 86M 84.5
ViL-B 89M 83.3
Table 9: Fine-tuning ImageNet-1K pre-trained models by using a 50% overlap when creating patches. This setting increases the number of patch tokens from 196 to 729. ViL easily outperforms Vim but falls short of ViTs on larger models. Note that ViL FLOPS are calculated using the parallel form of the mLSTM, which scales quadratically with the sequence length. Using the chunkwise form instead of the parallel form can reduce FLOPS for longer sequences (see Section C.2).
B.4 VTAB-1K Individual Dataset Results
Table 10 presents accuracies for each individual dataset of the VTAB-1K [64] benchmark.
Natural Specialized Structured
Cifar100
Caltech101
DTD
Flower102
Pets
SVHN
Sun397
Camelyon
EuroSAT
Resisc45
Retinopathy
Clevr-Count
Clevr-Dist
DMLab
KITTI-Dist
dSpr-Loc
dSpr-Ori
sNORB-Azim
sNORB-Ele
DeiT-T [51] 47.7 86.4 63.7 85.6 87.0 78.4 35.3 83.0 93.4 80.9 70.7 71.7 60.3 43.1 78.5 67.9 41.6 30.6 32.7
DeiT-III-T (reimpl.) 52.3 90.1 62.7 88.8 87.5 83.7 37.9 83.2 93.1 81.1 72.9 76.6 60.8 44.9 79.1 67.5 48.1 31.0 33.3
Vim-T [67] 46.7 86.3 60.7 84.0 88.8 76.1 33.7 82.2 92.9 75.2 72.6 59.8 49.9 39.3 78.2 51.2 43.9 26.9 27.2
ViL-T 54.2 90.2 67.4 90.7 89.9 81.6 41.1 83.4 94.2 82.7 73.1 80.7 61.8 49.4 81.3 57.8 51.8 31.4 34.8
DeiT-S [51] 57.0 88.9 68.2 90.9 90.8 75.4 42.1 83.3 94.0 83.8 74.0 74.6 58.3 45.6 78.2 61.9 47.9 27.1 31.9
DeiT-III-S [54] 58.8 88.6 67.5 90.9 91.7 84.4 43.3 84.4 92.6 82.5 73.5 76.5 57.9 46.2 78.9 58.3 49.7 23.7 27.5
Vim-S [67] 53.0 87.2 64.1 86.8 90.3 65.8 39.7 82.4 93.4 78.0 73.1 63.1 53.2 42.3 78.2 54.1 47.6 27.1 29.3
ViL-S 61.4 89.6 69.2 92.8 91.7 78.7 43.8 85.5 93.9 84.4 73.5 84.0 63.4 51.3 83.3 61.0 55.4 32.4 35.5
DeiT-B [51] 61.8 89.8 67.5 93.7 92.6 84.4 45.6 85.3 95.1 86.3 74.2 77.7 59.9 47.2 81.7 61.7 51.4 30.0 36.2
DeiT-III-B [54] 62.9 89.6 69.6 93.7 93.2 87.0 47.1 85.8 94.1 85.6 73.7 80.5 61.4 48.4 80.9 64.4 55.1 30.2 31.8
ViL-B 64.3 90.0 71.1 93.4 91.4 79.6 46.6 84.7 94.3 85.4 74.4 83.7 62.1 52.7 81.0 63.1 57.6 32.6 39.9
Table 10: Results on all datasets of the VTAB-1K [64] benchmark.
B.5 Robustness
Table 11 presents robustness and OOD evaluations of ImageNet-1K pre-trained classifiers.
Model IN-C ($\downarrow$) IN-A ($\uparrow$) IN-R ($\uparrow$) Sketch ($\uparrow$) Validation ($\uparrow$)
DeiT-T [51] 69.7 7.6 32.7 19.9 72.2
DeiT-III-T 65.0 11.7 39.4 27.4 76.2
Vim-T [67] 61.8 9.6 38.8 26.9 76.1
ViL-T 59.6 15.2 42.2 30.0 78.3
DeiT-S [51] 54.4 19.6 41.9 29.1 79.8
DeiT-III-S [54] 50.1 23.2 46.6 35.4 81.4
Vim-S [67] 51.5 19.7 44.8 32.5 80.5
ViL-S 50.6 23.8 47.9 35.2 81.5
DeiT-B [51] 48.6 27.9 44.6 32.0 81.8
DeiT-III-B [54] 42.7 36.5 54.1 41.1 83.8
ViL-B 45.3 30.9 51.9 39.0 82.4
Table 11: Robustness and OOD evaluations on ImageNet-C(orruption) [29], ImageNet-A(dversarial) [31], ImageNet-R(endition) [30] and ImageNet-Sketch [59].. For ImageNet-C, we report the mean corruption error [29] with AlexNet [39] as baseline.
Appendix C Implementation Details
C.1 Hardware
@@
We estimate the total number of A100 GPU-hours used for this project to be 38K hours. This estimate includes initial exploration, method development, analysis and evaluations.
C.2 FLOPS Calculation
We calculate the FLOPS with the fvcore library [^1] https://github.com/facebookresearch/fvcore. For ViL, we report the FLOPS of the mLSTM parallel form. Given sequence length $L$ and hidden dimension $D$, it has complexity $\mathcal{O}(L\frac{L}{2}D)$ where the $\frac{L}{2}$ stems from the causal masking.
We report FLOPS for a complexity of $\mathcal{O}(L\frac{L}{2}D)$ even though our current implementation of the mLSTM has complexity $\mathcal{O}(L^{2}D)$ (i.e., entries of the $QK$ matrix that are later masked are still calculated) which we justify by the fact that FlashAttention-2 [14] is approximately 1.7x faster with a causal mask than without. Therefore, an optimized hardware implementation of the mLSTM could also omit the calculation of values that are not needed.
As Vim [67] does not report FLOPS and their model makes use of CUDA kernels (which are not counted as FLOPS by fvcore), we replace all calls to CUDA kernels with their reference PyTorch implementation and count the FLOPS with fvcore.
For the total pre-training compute in Figure 2, we consider an efficient implementation of stochastic depth [32] which omits the calculation of a dropped block instead of masking it. Therefore, we change the implementation of ViT [18] to use our efficient stochastic depth implementation. Vim does not use stochastic depth for training as their models are relatively small.
C.3 ViL Hyperparameters
Table 12 shows detailed hyperparameters used to train ViL models.
@@ C.3 ViL Hyperparameters
Epochs 800 (T), 400 (S/B) $\rightarrow$ 20 (T, S), 5 (B)
Batch size 2048 $\rightarrow$ 1024
@@
   Base Learning rate 1e-3 $\rightarrow$ 1e-5
@@
Precision mixed bfloat16 [35]
   Backend torch.autocast
Learning rate schedule cosine decay
   Warmup schedule linear
   Warmup epochs 5 $\rightarrow$ 5 (T, S), 1 (B)
@@
Table 12: Hyperparameters for training ViL on ImageNet-1K, inspired by DeiT-III [54]. We follow the best setting from DeiT-III [54] and pre-train on 192 resolution followed by a short fine-tuning on 224 resolution (indicated by $\rightarrow$).
C.4 Fine-tuning on VTAB-1K
For fine-tuning models on VTAB-1K we provide the hyperparameters in Table 13. We search for the best learning rate for each dataset by fine-tuning the model 25 times (5 learning rates with 5 seeds each) on the 800 training samples and evaluating them on the 200 validation samples. With the best learning rate, we then train each model 5 times on concatenation of training and validation split, evaluate on the test split and report the average accuracy.
Parameter Value
Epochs 50
Batch size 64
Seeds 5
Optimizer AdamW [42, 36]
   Learning rate [1e-3, 7.5e-4, 5.0e-4, 2.5e-4, 1.0e-4]
   Layer-wise lr decay [13] 0.65*
   Weight decay 0.05
   Momentum $\beta_{1}=0.9,\beta_{2}=0.999$
Learning rate schedule linear warmup $\rightarrow$ cosine decay
   Warmup epochs 5
Precision mixed bfloat16 [35]
   Backend torch.autocast
Data Augmentation
   Resize
    interpolation bicubic
    size 224x224
   Normalize ImageNet-1K statistics
Table 13: Hyperparameters for fine-tuning on VTAB-1K. *For Vim and ViL we group two consecutive blocks for the layer-wise lr decay similar to how ViT considers a pair of attention and MLP block as a single “layer” for the decay.
C.5 ADE20K Semantic Segmentation Fine-tuning
We fine-tune models on ADE20K [66] using an UperNet [61] head. We follow common practices and fine-tune on 512x512 resolution, where we interpolate the absolute positional embedding from 224x224 to 512x512. For ViTs, we add relative position biases to the attention layers (initialized to 0) [27]. Table 14 lists detailed hyperparameters.
Parameter Value
Updates 160K
Batch size 16
UperNet
   Auxiliary
    Weight 0.4
    Input Block 8*
    Dimension 192 (T), 384 (S, B)
   Decoder
    Weight 1.0
    Input Blocks [4, 6, 8, 12]*
    Dimension 192 (T), 384 (S, B)
Stochastic depth [32]
   Peak rate 0 (T), 0.05 (S), 0.1 (B)
   Layer-wise Decay
Optimizer AdamW [42, 36]
   Learning rate 5e-4
   Linear LR Scaling Divisor [22] 16
   Layer-wise lr decay [13] 0.65*
   Weight decay 0.05
   Momentum $\beta_{1}=0.9,\beta_{2}=0.999$
Learning rate schedule linear warmup $\rightarrow$ cosine decay
   Warmup updates 1500
Precision mixed float16
   Backend torch.autocast
Train Data Augmentation
   RandomResize
    interpolation bicubic
   RandomCrop
    size 512x512
   RandomHorizontalFlip
   ColorJitter 0.5
    brightness 0.5
    contrast 0.5
    saturation 0.5
    hue 0.25
   Normalize ImageNet-1K statistics
Evaluation
   Stride 341
   Multi-scale
    scale factors [0.75, 1.0, 1.25, 1.5, 1.75]
    flip [True, False]
Table 14: Hyperparameters for fine-tuning on VTAB-1K. *For ViL we group two consecutive blocks into one similar to how a ViT block consists of a pair of attention and MLP block.
C.6 DeiT-III Reimplementation Hyperparameters
@@
Batch size 2048 $\rightarrow$ 1024
Model
   Patch size 16x16
   Latent dimension 192
   Depth 12
   Pooling [CLS]
@@
   Base Learning rate 1e-3 $\rightarrow$ 1e-5
@@
Precision mixed bfloat16 [35]
   Backend torch.autocast
Learning rate schedule cosine decay
   Warmup schedule linear